OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y > -x + 3? (0, 3) (0, 0) (1, -9) (2, 4)

OpenStudy (ajprincess):

Plug in the values and check which of the options satisfy the inequalty. let me show one. (0,3) x=0 y=3 -x+3=0+3 =3 y=3 but the y>-x+3 here it is equal. so (0,3) is not the answer. similarly can u check other options

OpenStudy (anonymous):

Is it (2,4)

OpenStudy (ajprincess):

yup:)